Using synthetic division to divide x^3-8x+7 by x-2 quotient and remainder
Answer by ewatrrr(24785)

 
Hi
Using synthetic division to divide x^3-8x+7 by x-2
2 1 0 -8 7
2 4 -8
1 2 -4 -1
(x^3-8x+7)/(x-2) = x^2 + 2x - 4 , R = -1/(x-2)